Assume the following information:

Amount Per Unit

Sales$300,000   $40   

Variable expenses 120,000    16   

Contribution margin 180,000   $24   

Fixed expenses 79,000        

Net operating income$101,000        

If the selling price per unit increases by 10% and unit sales drop by 5%, then the best of estimate of the new net operating income is:
Answer #1

Current units sold=($300,000/40)=7500 units

Hence new units sold=7500*(1-0.05)=7125 units

Sales(7125*(40*1.1)) 313500
Variable expenses(7125*16) 114,000
Contribution margin 199500
Fixed expenses 79000
Net operating income $120500

NOTE:Total fixed costs and variable cost per unit do not change with change in units